40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 3 18:42:43 2014, Jenne, Update, LSC, Green PDH box boosts BoostCartoon.pdf
    Reply  Thu Sep 4 00:56:44 2014, Jenne, Update, LSC, Green PDH box boosts TFstage_newDesign_3Sept2014.pngBoostNoBoost_newDesign_3Sept2014.pngBoostOnly_newDesign_3Sept2014.pngD0901351-v1_3Sept2014.pdf
       Reply  Thu Sep 4 03:12:55 2014, ericq, Update, LSC, Green PDH box boosts boostedTF.pdfgain.jpg
       Reply  Thu Sep 4 16:45:10 2014, Jenne, Update, LSC, Green PDH box boosts IMG_1712.JPGIMG_1714.JPG
          Reply  Fri Sep 5 05:32:57 2014, Jenne, Update, LSC, Green PDH box boosts TRX_TRY_ALSvsIR_4Sept2014.pdf
             Reply  Fri Sep 5 21:13:57 2014, Jenne, Update, LSC, Green PDH out of loop ALS_XY_outOfLoop_5Sept2014.pdf
Message ID: 10446     Entry time: Wed Sep 3 18:42:43 2014     Reply to this: 10448
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Green PDH box boosts 

From EricQ's simulations reported in elog 10390, we want to transition from ALS comm to DC transmission signals around 500 pm.  However, around 100 pm, the DC transmission signals have a sign flip, so we don't want to have the ALS swing that close to the CARM resonance.  So.  We want to be at about 500 pm, and not touch 100 pm.  So, we don't want our peak ALS motion to go beyond ~400 pm.  Which means that we need to have less than about 40 pm in-loop RMS, to avoid hitting 400 pm.  This is an ALS requirement, but since the analog PDH box is what forces the end laser to follow the arm cavity, and thus give us information about the arm length fluctuations, the PDH residual noise is part of our sensor noise for the full ALS.  So, we need to have the PDH in-loop RMS be less than 40 pm, integrated from a few kHz down to at least 30 mHz. Recall that above the ALS UGF (of about 200 Hz), the sensor noise will be suppressed by 1/f, so we should take that into account when we are looking at the PDH error signal, before we calculate the RMS motion.

Q also measured the in-loop error signal with the current Yend PDH box in elog 10430, and it looks like most of the RMS is coming from a few hundred Hz.  I designed a hack to the PDH board boost that has a zero at about 2kHz, and a gain of 30 at DC, so that we will win by squishing all that RMS.  Also, it shouldn't be too aggressive, so we should be able to leave it on all the time, and still acquire lock of the green laser to the arm, without having to do triggering.

The board schematic is at DCC D1400294.  The boost is also called the "integrator stage", although it will no longer be a simple integrator.

EDIT, JCD:  This cartoon is not correct for the non-boosted state, doesn't include effect of R16.

BoostCartoon.pdf

ELOG V3.1.3-